# Rotary Position Embedding for Vision Transformer

### v1

@@ Abstract
Keywords:
Relative position embedding Vision transformers Multi-resolution inference
@@ 1 Introduction
Transformers [34] have become popular in neural architecture due to their strong performance across various tasks in language and computer vision domains [5, 6]. The transformer treats input data as a sequence of tokens. The tokens equally interact with others through a self-attention mechanism [34]. Since the self-attention mechanism is independent of the token index or positions (i.e., permutation invariance), the transformer requires additional position information, usually injected by position embedding [34, 27, 5, 23]. The position embeddings give the position information to input tokens with specific embedding designed for the transformer. They uniquely differentiate tokens based on their locations rather than their contents. Thus, the position information of self-attention heavily depends on the position embedding, which is a crucial component in designing transformer architectures.
@@
This paper aims to improve position embedding for vision transformers by applying an extended Rotary Position Embedding (RoPE) [29]. RoPE is a relative position embedding that is specially designed for extrapolation in language domains. Despite the remarkable success of RoPE in Large-Language Models [33, 26, 12], its effectiveness in vision tasks has not been validated due to limited investigation. In this paper, we provide a comprehensive investigation of RoPE for transformers in vision recognition tasks. Our investigation starts with 1D to 2D expansion of RoPE to cope with images rather than original language inputs. Although 2D RoPE using axial frequencies was used in pioneer works [18, 7, 19], we argue that it lacks the ability to handle diagonal directions, which are preferred in convolution networks by the square kernel. To cope with the diagonal direction of RoPE, we propose to use mixed axis frequencies for 2D RoPE, named RoPE-Mixed. Since RoPE-Mixed uses frequencies for both axes as learnable parameters of the network, it effectively handles diagonal direction and is more suitable for ViT’s attention than Axial 2D RoPE.
In experiments, we apply variants of 2D RoPE to representative transformer architectures, ViT and Swin Transformer, and validate the effects of 2D RoPE in various tasks, including multi-resolution classification on ImageNet-1k [4], object detection on MS-COCO [16], and semantic segmentation on ADE20k [40, 41]. The results show that 2D RoPE is a beneficial option as position embedding for transformers with an impressive performance improvement on high-resolution images, i.e., extrapolation of images. We believe our study demonstrates the impact of 2D RoPE in vision domains and contributes to future research by suggesting a beneficial option in position embedding for vision transformers.
@@ 2 Related Works
Diverged from Convolutional Neural Networks (ConvNets) [8], ViT [6] requires a transformation in position embedding for multi-resolution inference. Some studies investigated a multi-resolution inference method for ViT to improve performance in downstream tasks. CAPE [15] analyzes ViT’s position embedding in resolution changes and finds that augmenting position embedding improves the multi-resolution performance of ViT. Thus, they propose a training recipe with continuous augmenting of position embedding (CAPE). ResFormer [30] shows that relative position embedding based on depth-wise convolution layer is beneficial for multi-resolution inference. Using this property, the study proposes an improved ViT architecture with global and local depth-wise conv embedding. It substantially improves multi-resolution performance with multi-resolution self-distillation learning recipes. In contrast to conventional multi-resolution, FlexiViT [1] proposes a ViT with flexible patch sizes that can replace multi-resolution inference. In FlexiViT, ViT increases the patch size instead of increasing input resolution. By training with a multi-patch-size training scheme and distillation using ViT-B/8 [28], FlexiViT exhibits remarkable performance for various patch-size, which corresponds to multi-resolution in computation cost aspect.
@@ 3.1.1 Absolute Positional Embedding (APE)[34, 5, 6]
is the most common position embedding for Vision Transformer (ViT). APE is generally added to the feature just after the patchification layer computes tokens from $16\times 16$ or $32\times 32$ patch images. For patch tokens $\mathbf{x}_{0}\in\mathbb{R}^{N\times d}$, APE $\mathbf{E}_{APE}\in\mathbb{R}^{N\times d}$ gives the position information for each token by addition:
@@
$\mathbf{E}_{APE}(\mathbf{p}_{n},4t)=\mathrm{sin}(p_{n}^{x}/10^{4t/\lfloor\frac{d}{4}\rfloor}),\,\mathbf{E}_{APE}(\mathbf{p}_{n},4t+1)=\mathrm{cos}(p_{n}^{x}/10^{4t/\lfloor\frac{d}{4}\rfloor})$ (2)
@@ 3.1.2 Relative Position Bias (RPB)
[23, 17] is a popular way to inject relative distances to ViT architectures. APE is not suitable for handling tokens based on their relative positions, as it relies solely on absolute positions in the image $\mathbf{p}_{n}=(p^{x}_{n},p^{y}_{n})$. It is necessary to use different types of position embedding that utilize relative positions $\tilde{\mathbf{p}}_{nm}=(\tilde{p}^{x}_{nm},\tilde{p}^{y}_{nm})=(p^{x}_{n}-p^{x}_{m},p^{y}_{n}-p^{y}_{m})$. RPB is widely used relative position embedding for ViT. In contrast to learnable APE, which has learnable parameters for each absolute position, RPB uses learnable parameters for each relative position. i.e., Relative Position Bias (RPB) table $T$ is defined as learnable parameters for every possible relative position:
@@
While APE is added to network features, RPB is directly applied to the attention matrix of at every self-attention layer since it is the only position that can handle relative relations in transformer architecture. The attention matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ with the query and key of a head denoted by $\mathbf{q},\mathbf{k}\in\mathbb{R}^{N\times d_{head}}$, is calculated
@@ 3.1.3 Rotary Position Embedding (RoPE)
where $\mathrm{Re}[\cdot]$ denotes real part of complex number and $*{}^{*}$ means complex conjugates. By multiplying complex rotation $e^{i\theta n},e^{i\theta m}$ depending on token position $(n,m)$, RoPE injects relative positions $(n-m)$ to the attention matrix in rotation form. In practical implementation, RoPE converts $\mathbf{q}_{n},\mathbf{k}_{m}\in\mathbb{R}^{1\times d_{head}}$ to complex vector $\mathbf{\bar{q}}_{n},\mathbf{\bar{k}}_{m}\in\mathbb{C}^{1\times(d_{head}/2)}$ by considers ($2t$)-th dim as real part and ($12t+$ 1)-th dim as imaginary part. It produces the same attention value as $\mathbf{q}_{n}\mathbf{k}_{m}^{T}=\mathrm{Re}[\mathbf{\bar{q}}_{n}\mathbf{\bar{k}}_{m}^{*}]$ but reduces computational wastes. Also, RoPE utilizes multiple frequencies $\theta_{t}$ using channel dimensions of key and query as
@@ 3.2 RoPE for 2D images
RoPE exhibits remarkable performance in the language domain. However, only a few studies have been conducted on using RoPE in the vision domain with 2D input, as it was designed solely for 1D input. This section introduces feasible implementations of 2D RoPE for image inputs: axial and learnable frequency.
@@
Note that $\theta_{t}$ for vision is often larger than that of language, and the number of frequencies is halved to cover both (x, y) dimensions with $d_{head}$ as well. This axial frequency has been used in a few pioneer works [7, 18, 19] to further improve the performance of a new ViT architecture.
@@
By using two frequencies for each axis, RoPE enables to handle the diagonal axis. The RoPE attention matrix in Eq. 8 is changed by mixed frequency as
@@
This formulation is identical to the axial frequency implementation as $\theta^{x}_{t}$ or $\theta^{y}_{t}$ goes to zero. Thus, mixed frequency RoPE is a generalized version of axial frequency RoPE. Different from fixed frequencies in language RoPE and axial frequency, we let the network learn frequencies $(\theta^{x}_{t},\theta^{y}_{t})$ for $t\in\{0,1,...,d_{head}/2\}$ as learnable parameters. Our mixed learnable frequency implementation enables diagonal direction handling to RoPE and makes RoPE learnable, like conventional positional embedding in the vision domain.Likeo RPB, we use separate sets of learnable frequencies for each head and every self-attention layer. It produces $d$ learnable parameters per self-attention layer. However, it is negligible since it requires only $\sim$0.01of network parameters in ViT-B.
@@ 3.3 Discussion
This section introduces the interesting features of RoPE for ViT, offering insights into its functionality from different perspectives.
@@
Vision models use diverse image resolutions depending on the goal of target tasks. For example, image classification uses $224\times 224$ as the standard resolution for comparison but utilizes small resolutions [32, 35] for training efficiency and enlarges resolutions to boost the performance additionally. Furthermore, object detection and segmentation prefer larger resolutions to capture small objects. Thus, transformers for vision should support resolution changes, which is linked to the necessity of resolution change in position embedding.
To address different resolutions, APE uses spatial resize with 2D interpolation. Similar to image resize function, $E_{APE}\in\mathbb{R}^{(H\times W)\times d}$ is fitted to different token size $E^{\prime}_{APE}=\mathrm{Resize}(E_{APE})\in\mathbb{R}^{(H^{\prime}\times W^{\prime})\times d}$ based on 2D interpolation such as bilinear or bicubic. In the case of RPB, the RPB table in Eq. 3 needs to be extended for large resolution. RPB applies zero-padding for large resolutions, which is the same as not using position embedding for extended resolutions. RoPE makes an extended position embedding based on sinusoidal function for large resolution. Different from zero-padding in RPB, the rotation matrix in Eq. 12 and Eq. 14 can produce values for extended positions since it is based on periodic functions, which has proven its effectiveness for extrapolation [33, 26]. We expect that the advantage of RoPE in extrapolation will also be effective for multi-resolution benchmark in §4.1 and dense prediction tasks in §4.2 and §4.3.
@@
Figure 1: Attention distances of ViT-B for APE/RoPE-Mixed. We measure the average distance of attention interaction by computing the distance between query-key tokens from attention probabilities. We average the distance across the validation set.
Figure 2: Entropy of attention in ViT-B with APE or RoPE-Mixed. Entropy of attention probability is measured for every self-attention of ViT-B. A high entropy value indicates that a large number of tokens are involved in the attention interaction.
We analyze the attention matrix of RoPE ViT-B compared to the original ViT with APE. Following attention analysis in previous literature [9, 20], we measure attention distances and entropy on ImageNet-1k validation set with various resolutions. Attention distance refers to the average spatial distance involved in attention interaction. The spatial distance between query-key is weighted and summed according to the attention probabilities. Note that we set the distance from the spatial token to the class token to zero for simplicity. Attention entropy represents the entropy values of attention probabilities, indicating probabilistic sharpness of attention. The measure is conducted for every self-attention layer except the last one since the last attention is only used for a class token. We use weights pretrained on ImageNet-1k in §4.1 and report average values across all tokens, attention heads, and validation samples.
The averaged attention distances for various resolutions are shown in Fig. 1. In training resolution $224\times 224$, RoPE-Mixed increases attention distance at the middle layers but decreases it in the second and later layers. In other resolutions, the pattern is similar, but the difference is more significant than the training resolution. In short, RoPE-Mixed increases attention distance at the middle layers compared to APE, which becomes substantial at resolution changes. The entropy results are reported in Fig. 2 on multiple resolutions. Interestingly, the pattern is similar to attention distance. The entropy of RoPE-Mixed is larger than APE at middle layers. These analysis results imply that RoPE-Mixed makes attention interact with long-range (attention distance) and various tokens (entropy). We speculate that these differences in attention contributed to the performance improvement of RoPE observed in §4.
@@ 4 Experiments
We apply 2D RoPE to two representative vision transformer architectures: ViT [6] and Swin Transformer [17]. Note that ViT uses APE, whereas Swin Transformer uses RPB. Thus, our experiment can verify the performance of RoPE when it replaces APE or RPB. RoPE in ViT and Swin Transformer is validated for vision recognition tasks, including multi-resolution classification (§4.1) on ImageNet-1k [4], object detection (§4.2) on MS-COCO [16], and semantic segmentation (§4.3) on ADE20k [40, 41]. On various tasks, we compare the conventional position embeddings (APE, RPB) with two variants of 2D RoPE RoPE-Axial (Eq. 12) and RoPE-Mixed (Eq. 14). Our experiments will exhibit the remarkable performance of 2D RoPE across all tasks, particularly with a significant margin in extrapolation.
@@ 4.1 Multi-resolution classification
Robustness on multi-resolution inputs is an essential factor of ViT performance, as it is closely related to their downstream performance in dense prediction tasks. In language models [33, 26, 12], RoPE exhibited strong extrapolation performance, i.e., text sequence longer than training samples. 2D RoPE might be suitable for large-resolution images, leveraging its extrapolation capabilities. We train ViTs and Swin Transformers on ImageNet-1k [4] training set with high-performance training recipes [32, 17]. We report the accuracy on the ImageNet-1k validation set as varying image sizes. Note that we use the ImageNet-1k standard image resolution $224\times 224$ for training. Thus, a resolution larger than 224 is considered as extrapolation.

### v2

@@ 1 Introduction
Transformers [34] have become popular due to their strong performance across various tasks in language and computer vision domains [5, 6]. The transformer treats input data as a sequence of tokens. The tokens equally interact with others through a self-attention mechanism [34]. Since the self-attention mechanism is independent of the token index or positions (i.e., permutation invariance), the transformer requires additional position information, usually injected by position embedding [34, 27, 5, 23]. The position embeddings give the position information to input tokens with specific embedding designed for the transformer. They uniquely differentiate tokens based on their locations rather than their contents. Thus, the position information of self-attention heavily depends on the position embedding, which is a crucial component in designing transformer architectures.
@@
This paper aims to improve position embedding for vision transformers by applying an extended Rotary Position Embedding (RoPE) [29]. RoPE is a relative position embedding that is specially designed for extrapolation in language domains. Despite the remarkable success of RoPE in Large-Language Models [33, 26, 12], its effectiveness in vision tasks has not been validated due to limited investigation. In this paper, we provide a comprehensive investigation of RoPE for transformers in vision recognition tasks. Our investigation starts with 1D to 2D expansion of RoPE to cope with images rather than original language inputs. Although 2D RoPE using axial frequencies was used in pioneer works [18, 7, 19], we argue that it lacks the ability to handle diagonal directions, which are preferred in convolution networks by the square kernel. To cope with the diagonal direction of RoPE, we propose to use mixed axis frequencies for 2D RoPE, named RoPE-Mixed. Since RoPE-Mixed uses frequencies for both axes as learnable network parameters, it effectively handles diagonal direction and is more suitable for ViT’s attention than Axial 2D RoPE.
In experiments, we apply variants of 2D RoPE to representative transformer architectures, ViT and Swin Transformer, and validate the effects of 2D RoPE in various tasks, including multi-resolution classification on ImageNet-1k [4], object detection on MS-COCO [16], and semantic segmentation on ADE20k [40, 41]. The results show that 2D RoPE is a beneficial option for position embedding in transformers with impressive performance improvements on high-resolution images, i.e., extrapolation of images. We believe our study demonstrates the significant impact of 2D RoPE in vision domains and contributes to future research by suggesting a beneficial option in position embedding for vision transformers.
@@ 2 Related Works
Unlike ConvNets [8], ViT [6] requires a transformation in position embedding for multi-resolution inference. Some studies investigated a multi-resolution inference method for ViT. CAPE [15] analyzes ViT’s position embedding in resolution changes and finds that augmenting position embedding improves the multi-resolution performance of ViT. Thus, they propose a new training recipe that includes continuous augmenting of position embedding (CAPE). ResFormer [30] shows that relative position embedding based on depth-wise convolution layer benefits multi-resolution inference. Using this property, the study proposes an improved ViT architecture with global and local depth-wise conv embedding. It substantially improves multi-resolution performance with multi-resolution self-distillation learning recipes. In contrast to conventional multi-resolution, FlexiViT [1] proposes a ViT with flexible patch sizes that can replace multi-resolution inference. In FlexiViT, ViT increases the patch size instead of increasing input resolution. By training with a multi-patch-size training scheme and distillation using ViT-B/8 [28], FlexiViT exhibits remarkable performance for various patch-size, which corresponds to multi-resolution in computation cost aspect.
@@ 3.1.1 Absolute Positional Embedding (APE)[34, 5, 6]
is the most common position embedding for Vision Transformer (ViT). APE is generally added to the feature right after the patchification layer computes tokens from $16\times 16$ or $32\times 32$ patch images. For patch tokens $\mathbf{x}_{0}\in\mathbb{R}^{N\times d}$, APE $\mathbf{E}_{APE}\in\mathbb{R}^{N\times d}$ gives the position information for each token by addition:
@@
$\mathbf{E}_{APE}(\mathbf{p}_{n},4t)=\mathrm{sin}(p_{n}^{x}/10^{4t/\lfloor\frac{d}{4}\rfloor}),\,\mathbf{E}_{APE}(\mathbf{p}_{n},4t+1)=\mathrm{cos}(p_{n}^{x}/10^{4t/\lfloor\frac{d}{4}\rfloor}),$ (2)
@@ 3.1.2 Relative Position Bias (RPB)
[23, 17] is a popular way to inject relative distances to the ViT architectures. APE is not suitable for handling tokens based on their relative positions, as it relies solely on absolute positions in the image $\mathbf{p}_{n}=(p^{x}_{n},p^{y}_{n})$. It is necessary to use different types of position embedding that utilize relative positions $\tilde{\mathbf{p}}_{nm}=(\tilde{p}^{x}_{nm},\tilde{p}^{y}_{nm})=(p^{x}_{n}-p^{x}_{m},p^{y}_{n}-p^{y}_{m})$. RPB is widely used relative position embedding for ViT. In contrast to learnable APE, which has learnable parameters for each absolute position, RPB uses learnable parameters for each relative position. i.e., Relative Position Bias (RPB) table $T$ is defined as learnable parameters for every possible relative position:
@@
While APE is added to network features, RPB is directly applied to the attention matrix of every self-attention layer since it is the only position that can handle relative relations in transformer architecture. The attention matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ with the query and key of a head denoted by $\mathbf{q},\mathbf{k}\in\mathbb{R}^{N\times d_{head}}$, is calculated
@@ 3.1.3 Rotary Position Embedding (RoPE)
where $\mathrm{Re}[\cdot]$ denotes real part of complex number and means complex conjugates. By multiplying complex rotation $e^{i\theta n},e^{i\theta m}$ depending on token position $(n,m)$, RoPE injects relative positions $(n-m)$ to the attention matrix in rotation form. In practical implementation, RoPE converts $\mathbf{q}_{n},\mathbf{k}_{m}\in\mathbb{R}^{1\times d_{head}}$ to complex vector $\mathbf{\bar{q}}_{n},\mathbf{\bar{k}}_{m}\in\mathbb{C}^{1\times(d_{head}/2)}$ by considers ($2t$)-th dim as real part and ($12t+$ 1)-th dim as imaginary part. It produces the same attention value as $\mathbf{q}_{n}\mathbf{k}_{m}^{T}=\mathrm{Re}[\mathbf{\bar{q}}_{n}\mathbf{\bar{k}}_{m}^{*}]$ but reduces computational wastes. Also, RoPE utilizes multiple frequencies $\theta_{t}$ using channel dimensions of key and query as
@@ 3.2 RoPE for 2D images
RoPE exhibits remarkable performance in the language domain. However, only a few studies have explored using RoPE in the vision domain with 2D input, as it was designed solely for 1D input. This section introduces feasible implementations of 2D RoPE for input images: axial and learnable frequency.
@@
Note that $\theta_{t}$ for vision is often larger than that of language, and the number of frequencies is halved to cover both (x, y) dimensions with $d_{head}$ as well. This axial frequency has been used in a few pioneering works [7, 18, 19] to further improve the performance of a new ViT architecture.
@@
By using two frequencies for each axis, RoPE allows handling of the diagonal axis. The RoPE attention matrix in Eq. 8 is changed by mixed frequency as
@@
This formulation is identical to the axial frequency implementation as $\theta^{x}_{t}$ or $\theta^{y}_{t}$ goes to zero. Thus, mixed frequency RoPE is a generalized version of axial frequency RoPE. Different from fixed frequencies in language RoPE and axial frequency, we let the network learn frequencies $(\theta^{x}_{t},\theta^{y}_{t})$ for $t\in\{0,1,...,d_{head}/2\}$ as learnable parameters. Our mixed learnable frequency implementation enables diagonal direction handling to RoPE and makes RoPE learnable, like conventional positional embedding in the vision domain. Like RPB, we use separate sets of learnable frequencies for each head and every self-attention layer. It produces $d$ learnable parameters per self-attention layer. However, it is negligible since it requires only $\sim$0.01of network parameters in ViT-B.
@@ 3.3 Discussion
3.3.1 2D Fourier analysis.
We design a 2D Fourier analysis to demonstrate the representational difference between RoPE-Axial and RoPE-Mixed. When all 2D frequencies are utilized, a 2D Fast Fourier Transform (FFT) followed by an inverse Fast Fourier Transform (iFFT) perfectly reconstructs the input. However, the number of RoPE frequencies is limited to $\frac{d_{head}}{2}$ as in Eq. 9. $\frac{d_{head}}{2}$ ($=32$ for ViT-B) frequencies are insufficient to cover all 2D frequencies, resulting in imperfect reconstructions. This imperfect reconstruction reflects the expressiveness and representation pattern of the frequencies. In Fig. 1, we compare 2D FFT-iFFT results of RoPE-Axial and RoPE-Mixed frequencies. Note that we use RoPE-Mixed frequencies from ViT-B trained on ImageNet-1k. The results show a significant difference: Axial frequencies exhibit artifacts along axial lines, impairing precise positional representation, whereas Mixed frequencies utilize diverse 2D frequencies to produce sharper locations. Thus, we claim that the mixed frequencies are necessary for precise localization in the attention, which explains why RoPE-Mixed performs better than RoPE-Axial in §4.
Figure 1: 2D Fourier reconstruction with RoPE frequencies. We perform a Fast Fourier Transform (FFT) followed by an inverse FFT with only RoPE frequencies to evaluate the representation capabilities of RoPE frequencies
@@
Vision models use diverse image resolutions depending on the goal of target tasks. For example, image classification uses $224\times 224$ as the standard resolution for comparison but utilizes small resolutions [32, 35] for training efficiency and enlarges resolutions to boost the performance additionally. Furthermore, object detection and segmentation prefer larger resolutions to capture small objects. Thus, transformers for vision should support resolution changes, which is linked to the necessity of resolution change in position embedding. RoPE makes an extended position embedding based on sinusoidal function for large resolution. Different from zero-padding in RPB, the rotation matrix in Eq. 12 and Eq. 14 can produce values for extended positions since it is based on periodic functions, which has proven its effectiveness for extrapolation [33, 26]. We expect that the advantage of RoPE in extrapolation will also be effective for multi-resolution benchmark in §4.1 and dense prediction tasks in §4.2 and §4.3.
@@
Figure 2: Attention distances of ViT-B for APE/RoPE. We measure the average distance of attention interaction by computing the distance between query-key tokens from attention probabilities. We average the distance across the validation set.
Figure 3: Entropy of attention in ViT-B with APE or RoPE. Entropy of attention probability is measured for every self-attention of ViT-B. A high entropy value indicates that a large number of tokens are involved in the attention interaction.
We analyze the attention matrix of RoPE ViT compared to the ViT with APE. Following attention analysis in literature [9, 20], we measure attention distances and entropy on the ImageNet-1k validation set with various resolutions. Attention distance refers to the average spatial distance involved in attention interaction. Attention entropy represents the entropy values of attention probabilities, indicating the sharpness of attention. The averaged attention distances are shown in Fig. 2. In training resolution $224\times 224$, RoPEs increase attention distance at the middle layers but decrease it in the second and later layers. In other resolutions, the pattern is similar, but the difference is more significant than the training resolution. In short, RoPEs increase attention distance at the middle layers, which becomes substantial at resolution changes. The entropy results are reported in Fig. 3. Interestingly, the pattern is similar to attention distance. Entropy of RoPE is larger than that of APE at the middle layers. These analysis results imply that RoPE makes attention interact with long-range (attention distance) and various tokens (entropy). We speculate that these differences in attention contributed to the performance improvement of RoPE observed in §4.
@@ 4 Experiments
We apply 2D RoPE to two representative ViT architectures: ViT [6] and Swin Transformer [17]. Note that ViT uses APE, whereas Swin Transformer uses RPB. Thus, our experiment can verify the performance of RoPE when it replaces APE or RPB. RoPE in ViT and Swin Transformer is validated for image recognition, including multi-resolution classification (§4.1) on ImageNet-1k [4], object detection (§4.2) on MS-COCO [16], and semantic segmentation (§4.3) on ADE20k [40, 41]. We compare the conventional position embeddings (APE, RPB) with two variants of 2D RoPE RoPE-Axial (Eq. 12) and RoPE-Mixed (Eq. 14). Our experiments will exhibit the remarkable performance of 2D RoPE across all tasks, particularly with a significant margin in extrapolation.
@@ 4.1 Multi-resolution classification
Robustness on multi-resolution inputs is an essential factor of ViT performance, as it is closely related to their downstream performance in dense prediction tasks. In language models [33, 26, 12], RoPE exhibited strong extrapolation performance, i.e., text sequence longer than training samples. 2D RoPE might be suitable for large-resolution images, leveraging its extrapolation capabilities. We train ViTs and Swin Transformers on ImageNet-1k [4] training set with high-performance training recipes [32, 17]. We report the accuracy on the ImageNet-1k validation set as varying image sizes. Note that we use the ImageNet-1k standard image resolution $224\times 224$ for training. Thus, a resolution larger than 224 can be considered as extrapolation.